A small ball is thrown vertically upwards from the top of a tower with an initial velocity of 20 m/s. If the ball took a total of 6 s to reach the ground level, determine the height of the tower.
1
Expert's answer
2020-07-02T17:15:20-0400

Split the time into several parts. First, calculate how long it will take to reach the highest point of trajectory if the body is launched vertically at 20 m/s:


"t_1=\frac{v}{g}."


Then the body falls from the maximum height "H" and this takes time "t_2=t-t_1":


"H=\frac{gt_2^2}{2}=\frac{g(t-t_1)^2}{2}=\frac{g(t-v/g)^2}{2}."

The height above the tower that the body reached is


"y=\frac{v^2}{2g}."

Thus, the height of the tower is


"h=H-y=\frac{g(t-v/g)^2}{2}-\frac{v^2}{2g},\\\space\\
h=\frac{9.8(6-20/9.8)^2}{2}-\frac{20^2}{2\cdot9.8}=56.4\text{ m}."